# Spin excitations of the Shastry-Sutherland model – altermagnetism and proximate deconfined quantum criticality

### v1

@@ III Numerical results on the spin excitation spectra
Figure S-2: Total spin DSF ($\mathcal{S}^{xx+yy+zz}$) in the AFM phase of the SS model calculated by iPEPS method with $D=3$ (in (a) and (b)) and $D=2$ (in (c) and (d)), at $J/J^{\prime}=0.9$ (in (a) and (c)) and $J/J^{\prime}=1.1$ (in (b) and (d)).
@@
Figure S-3: Finite $D$ scaling of the spin gap of the Goldstone modes (at M point) in the AFM phase of the SS model at $J/J^{\prime}=1.1$. The dashed line shows a best fit to the gap, which vanishes in the large $D$ limit.
@@
Here we show numerical results on the spin excitation spectra in additional to those in the main text. The spectra illustrate the spin DSF In Fig. 2(b) we have shown the spectra in the AFM phase at $J/J^{\prime}=1.1$ calculated in the TN approach with bond dimension $D=3$. In Fig. S-2, we show calculated spectra for several other $J/J^{\prime}$ values and with different bond dimension $D$. One sees that the spectra show similar features with two split magnon bands, which are the Goldstone modes, at low energy and several longitudinal modes at high energy. Interestingly, the spin gap of the Higgs mode decreases with decreasing $J/J^{\prime}$ toward the plaquette-to-AFM transition, as discussed in the main text.
@@
As an effect of finite $D$, the Goldstone modes exhibit a finite gap. We have calculated the spin gaps with other $D$ values. As shown in Fig. S-3, the spin gap reduces with increasing $D$ and eventually vanishes in the large $D$ limit. This indicates that the gapless feature of Goldstone modes is captured by the tensor network calculation in this limit.
@@
Figure S-4: (a): Difference of the transverse DSF $\Delta\mathcal{S}=\mathcal{S}^{+-}(\mathbf{q},\omega)-\mathcal{S}^{-+}(\mathbf{q},\omega)$ in the AFM phase of the SS model at $J/J^{\prime}=1.1$ in the LSW theory. (b): Same as (a), but calculated by the tensor network method. In either case, the opposite sign of $\Delta\mathcal{S}$ reflects the opposite chirality of the magnon modes.
@@
To show that the chiral magnon excitations in the AFM phase of the SS model, we calculate the difference of the transverse component of the DSF $\Delta\mathcal{S}=\mathcal{S}^{+-}(\mathbf{q},\omega)-\mathcal{S}^{-+}(\mathbf{q},\omega)$. Note that $S_{i}^{+}S_{j}^{-}-S_{i}^{-}S_{j}^{+}\propto(\mathbf{S}_{i}\times\mathbf{S}_{j})_{z}$, so that the sign of $\Delta\mathcal{S}$ is able to detect the chirality of the spin excitations. In Fig. S-4 we compare results of $\Delta\mathcal{S}$ calculated by LSW theory and tensor network method. Both results show a finite splitting between the two Goldstone modes with opposite sign of $\Delta\mathcal{S}$. Also note that the magnon bands along $\Gamma$-M and $\Gamma$-M also have opposite sign. Interestingly, we find that the optical magnon bands also exhibit opposite chirality. These results verify the existence of chiral magnons in the AFM state, which is caused by the symmetry of altermagnetism.
@@
Figure S-5: (a): Density of states (DoS) with excitation energy at several $J/J^{\prime}$ values in the AFM phase. (b): Same as (a) but in the plaquette phase.
@@
Besides the collective magnon excitations and the longitudinal modes, the tensor calculation also resolves continuum excitations. The spectral weights of these excitations are much smaller than the collective modes but as shown in Fig. S-5, they cause a broad peak in the density of states (DoS) at high energy. In the AFM phase, this broad peak appears in energies above the Higgs gap (above $\sim 2J^{\prime}$ at $J/J^{\prime}=1.1$), and the peak energy decreases with varying $J/J^{\prime}$ toward the plaquette-to-AFM transition. This behavior is consistent with the emergence of deconfined excitations although it is hard to tell whether the continuum has a spinon or magnon origin. In the plaquette phase, we can also identify continuous excitations at energies above the triplet gap (above $\sim J^{\prime}$ at $J/J^{\prime}=0.7$) in the DoS plots (Fig. S-5(b)). The behavior of energies of these excitations is similar to that in the AFM phase.
@@
Figure S-6: (a): Field dependence of gaps of several low-energy excitations in the plaquette phase of the SS model, calculated by the TN method. The multiplets split into $S^{z}=+1,-1,0$ states. (b): Field dependence of low-energy excitations of a 4-spin cluster with nearest neighbor AFM interactions, calculated by the ED. The splitting indicates that the 7-fold multiplet contains two degenerate triplets and one singlet.
@@
To understand the singlet and the 7-fold multiplet modes above the lowest triplet mode in the plaquette phase resolved by the tensor network calculation in the main text, we apply a small magnetic field to split the multiplets. We also compare our tensor network results with those from exact diagonalization on a four-spin plaquette with nearest neighbor Heisenberg interaction. As shown in Fig. S-6(a), the 7-fold quasi-degenerate excitations splits under a small field. The gaps of three modes are unchanged under the field, while the gaps of two modes increase with the field, gaps of another two modes decrease with the field. Note that the slops of the gaps are the same for these modes, and they are also same to the slops of the split lowest triplet excitations. These indicate that the 7-fold multiplet actually consists of one singlet and two degenerate triplets. This conclusion is further confirmed by our exact diagonalization calculation shown in Fig. S-6(b). At zero field, there are indeed 7-fold degenerate multiplet excitations above the triplet ones. With applying a small magnetic field, The 7-fold multiplet splits into a pair of triplets and one singlet. Note that the singlet excitation in between the lowest triplet and the 7-fold multiplet seen in the tensor network result in Fig. 3 of the main text does not show up in our 4-spin cluster exact diagonalization calculation. It is then interpreted as an inter-plaquette bound state.
@@
Figure S-7: Evolution of gaps at the M point of the triplet mode (TM), Higgs mode (HM), and Goldstone modes (GM) with $J/J^{\prime}$ near the plaquette-to-AFM transition in the iPEPS calculation with $D=3$ (in (a)) and $D=5$ (in (b)). In either case we observe softening of gaps of both the triplet mode and Higgs mode when approaching the plaquette-to-AFM transition.
@@
In Fig. 4(a) of the main text we discuss the evolution of triplet mode, Higgs mode, and Goldstone modes with varying $J/J^{\prime}$. It is found that the gaps of both triplet mode and Higgs mode drop rapidly to small values close to zero, and become quasi-degenerate with the Goldstone modes when approaching the plaquette-to-AFM transition. Here in Fig. S-7 we show results with several other bond dimension $D$ values. We observe the same behavior that the gaps of both triplet mode and Higgs mode are softened near the transition. This behavior strongly suggests emergence of nearly deconfined excitations at the plaquette-to-AFM transition.

### v2

@@ Appendix C Numerical results on the spin excitation spectra
Figure 6: Total spin DSF ($\mathcal{S}^{xx+yy+zz}$) in the AFM phase of the SS model calculated by iPEPS method with $D=3$ (in (a) and (b)) and $D=2$ (in (c) and (d)), at $J/J^{\prime}=0.9$ (in (a) and (c)) and $J/J^{\prime}=1.1$ (in (b) and (d)).
@@
Figure 7: Finite $D$ scaling of the spin gap of the Goldstone modes (at M point) in the AFM phase of the SS model at $J/J^{\prime}=1.1$. The dashed line shows a best fit to the gap, which vanishes in the large $D$ limit.
Here we show numerical results on the spin excitation spectra in additional to those in the main text. The spectra illustrate the spin DSF
$\mathcal{S}^{\alpha\beta}(\mathbf{q},\omega)=\frac{1}{N}\sum_{i,j}\int dte^{i\mathbf{q}\cdot(\mathbf{r}_{i}-\mathbf{r}_{j})}e^{i\omega t}\langle S_{i}^{\alpha}(t)S_{j}^{\beta}(0)\rangle$ (24)
@@
with spin components $\alpha,\beta$. In Fig. 2(b) we have shown the spectra in the AFM phase at $J/J^{\prime}=1.1$ calculated in the tensor network approach with bond dimension $D=3$. In Fig. 6, we show calculated spectra for several other $J/J^{\prime}$ values and with different bond dimension $D$. One sees that the spectra show similar features with two split magnon bands, which are the Goldstone modes, at low energy and several longitudinal modes at high energy. Interestingly, the spin gap of the Higgs mode decreases with decreasing $J/J^{\prime}$ toward the PVBS-to-AFM transition, as discussed in the main text.
@@
As an effect of finite $D$, the Goldstone modes exhibit a finite gap. We have calculated the spin gaps with other $D$ values. As shown in Fig. 7, the spin gap reduces with increasing $D$ and eventually vanishes in the large $D$ limit. This indicates that the gapless feature of Goldstone modes is captured by the tensor network calculation in this limit.
@@
Figure 8: (a): Difference of the transverse DSF $\Delta\mathcal{S}=\mathcal{S}^{+-}(\mathbf{q},\omega)-\mathcal{S}^{-+}(\mathbf{q},\omega)$ in the AFM phase of the SS model at $J/J^{\prime}=1.1$ in the LSW theory. (b): Same as (a), but calculated by the tensor network method. In either case, the opposite sign of $\Delta\mathcal{S}$ reflects the opposite chirality of the magnon modes.
@@
To show that the chiral magnon excitations in the AFM phase of the SS model, we calculate the difference of the transverse component of the DSF $\Delta\mathcal{S}=\mathcal{S}^{+-}(\mathbf{q},\omega)-\mathcal{S}^{-+}(\mathbf{q},\omega)$. Note that $S_{i}^{+}S_{j}^{-}-S_{i}^{-}S_{j}^{+}\propto(\mathbf{S}_{i}\times\mathbf{S}_{j})_{z}$, so that the sign of $\Delta\mathcal{S}$ is able to detect the chirality of the spin excitations. In Fig. 8 we compare results of $\Delta\mathcal{S}$ calculated by LSW theory and tensor network method. Both results show a finite splitting between the two Goldstone modes with opposite sign of $\Delta\mathcal{S}$. Also note that the magnon bands along $\Gamma$-M and $\Gamma$-M also have opposite sign. Interestingly, we find that the optical magnon bands also exhibit opposite chirality. These results verify the existence of chiral magnons in the AFM state, which is caused by the symmetry of altermagnetism.
@@
Figure 9: (a): Density of states (DoS) with excitation energy at several $J/J^{\prime}$ values in the AFM phase. (b): Same as (a) but in the plaquette phase.
@@
Besides the collective magnon excitations and the longitudinal modes, the tensor calculation also resolves continuum excitations. The spectral weights of these excitations are much smaller than the collective modes but as shown in Fig. 9, they cause a broad peak in the density of states (DoS) at high energy. In the AFM phase, this broad peak appears in energies above the Higgs gap (above $\sim 2J^{\prime}$ at $J/J^{\prime}=1.1$), and the peak energy decreases with varying $J/J^{\prime}$ toward the PVBS-to-AFM transition. This behavior is consistent with the emergence of deconfined excitations although it is hard to tell whether the continuum has a spinon or magnon origin. In the plaquette phase, we can also identify continuous excitations at energies above the triplet gap (above $\sim J^{\prime}$ at $J/J^{\prime}=0.7$) in the DoS plots (Fig. 9(b)). The behavior of energies of these excitations is similar to that in the AFM phase.
@@
Figure 10: (a): Field dependence of gaps of several low-energy excitations in the plaquette phase of the SS model, calculated by the tensor network method. The multiplets split into $S^{z}=+1,-1,0$ states. (b): Field dependence of low-energy excitations of a 4-spin cluster with nearest neighbor AFM interactions, calculated by the ED. The splitting indicates that the 7-fold multiplet contains two degenerate triplets and one singlet.
@@
To understand the singlet and the 7-fold multiplet modes above the lowest triplet mode in the plaquette phase resolved by the tensor network calculation in the main text, we apply a small magnetic field to split the multiplets. We also compare our tensor network results with those from exact diagonalization on a four-spin plaquette with nearest neighbor Heisenberg interaction. As shown in Fig. 10(a), the 7-fold quasi-degenerate excitations splits under a small field. The gaps of three modes are unchanged under the field, while the gaps of two modes increase with the field, gaps of another two modes decrease with the field. Note that the slops of the gaps are the same for these modes, and they are also same to the slops of the split lowest triplet excitations. These indicate that the 7-fold multiplet actually consists of one singlet and two degenerate triplets. This conclusion is further confirmed by our exact diagonalization calculation shown in Fig. 10(b). At zero field, there are indeed 7-fold degenerate multiplet excitations above the triplet ones. With applying a small magnetic field, The 7-fold multiplet splits into a pair of triplets and one singlet. Note that the singlet excitation in between the lowest triplet and the 7-fold multiplet seen in the tensor network result in Fig. 3 of the main text does not show up in our 4-spin cluster exact diagonalization calculation. It is then interpreted as an inter-plaquette bound state.
@@
Figure 11: Evolution of gaps at the M point of the triplet (T) and singlet (S) modes in the PVBS phase, as well as the Higgs (H), magnon (M), and singlet (S) modes in the AFM phase, with $J/J^{\prime}$ near the PVBS-to-AFM transition in the iPEPS calculation with $D=3$ (in (a)) and $D=4$ (in (b)). In either case we observe softening of all collective modes when approaching the transition point.
@@
In Fig. 4(a) of the main text we discuss the evolution of various collective excitations with varying $J/J^{\prime}$. These include triplet and singlet modes in the PVBS phase, and the Higgs, magnon, and singlet modes in the AFM phase. It is found in Fig. 4(c) that the gaps of triplet and singlet modes in the PVBS phase, together with those of the Higgs and singlet modes in the AFM phase, drop rapidly to small values close to zero, and become quasi-degenerate with the magnon (Goldstone) modes when approaching the PVBS-to-AFM transition. Here in Fig. 11 we show results with smaller bond dimensions $D=3$ and $D=4$. We observe the same behavior that all these collective modes are softened near the transition. Note that the gap between the Higgs and magnon modes is still sizable at the transition point for $D=3$ and $D=4$. But as shown in Fig. 4(c) and (e), this gap drops rapidly with increasing $D$ and eventually closes in the large $D$ limit. This behavior strongly suggests the emergence of a DQCP at the PVBS-to-AFM transition.